### v1

@@ Abstract
Keywords:
Attack graphs Active Directory Stackelberg games Graph neural networks Evolutionary algorithms Adaptive cyber defence
@@ 1 Introduction
Enterprise networks are increasingly targeted by sophisticated adversaries who exploit configuration flaws and access dependencies to escalate privileges and compromise high-value assets [2, 11]. Defenders, meanwhile, must respond under uncertainty and with limited resources, making proactive, strategic countermeasures essential. These challenges are amplified in complex enterprise infrastructures, where hierarchical access and interdependent components expand the attack surface [8, 9].
A particularly critical target in this landscape is Microsoft Active Directory (AD), the central identity and access management system for Windows-based networks [4]. Over 90% of Fortune 1000 companies rely on AD to govern authentication, authorisation, and access control [10], making it a high-value target. According to Enterprise Management Associates (EMA 2021), more than half of surveyed organisations experienced AD-related breaches. These breaches often follow an “identity snowball” pattern, where attackers pivot from low-privilege accounts toward the highly privileged Domain Admin (DA) node via lateral movement [3]. AD environments can be represented as attack graphs, with nodes denoting entities (e.g., users, machines, groups) and edges representing access relationships. Tools such as BloodHound [1] are used to analyse AD graphs by identifying shortest paths to high-value targets, but they assume static attacker strategies (these strategies assume fixed, precomputed paths that remain unchanged despite defensive actions) and fail to model adaptive behaviour. In practice, attackers frequently adapt to blocked or failed paths by exploiting alternate paths, credentials, underused privileges, and overlooked misconfigurations. Prior work for defending AD graphs has explored edge-blocking techniques that selectively block high-impact edges. However, these methods often assume fixed attacker strategies and lack mechanisms to model dynamic adaptation, limiting their robustness and generalisability to real-world threats.
@@ 4.1 Graph Reduction via Fixed-Parameter Tractable Procedure
Condensing the AD Graph. We propose a Fixed-Parameter Tractable (FPT) procedure to preprocess the original AD graph into a smaller yet strategically equivalent representation. Intuitively, an AD graph can be viewed as a spanning tree augmented by $h$ feedback edges, where $h=m-(n-1)$ and $m$ and $n$ denote the number of edges and nodes, respectively. Within this graph, Splitting nodes ($t$) are nodes with multiple outgoing edges, representing critical decision points for the attacker, and Entry nodes ($s$) are nodes from which the attacker can initiate an attack. Let SplitSet represent the set of all splitting nodes and EntrySet represent the set of all entry nodes. Our FPT procedure leverages these sets to systematically condense the graph while preserving the structural properties critical to attacker–defender interactions. To facilitate this, we use the concept of Non-Splitting Paths [5]:
@@
[…]
@@ 4.3 Defender Strategy Optimisation via Evolutionary Diversity Optimisation (EDO)
[…]
@@
The newly generated individual is retained in the population if its fitness, defined as the attacker’s success rate under the corresponding blocking plan, as predicted by the GNN, falls within a tolerance window around the current best score. Specifically, it must lie within the range $[f^{*}-0.1,\ f^{*}+0.1]$, where $f^{*}$ is the best fitness value observed so far. This 0.1 margin ensures that near-optimal plans are preserved while maintaining population diversity, unless the new individual significantly improves overall diversity.
@@
$D(C(bw)\backslash{p_{i}})=C(bw)-p_{i}$
$D(C(bw)\backslash{p_{i}})=\Big{(}c(bw_{1})-(bw_{1};i),...,c(bw_{|BW|})-(bw_{|BW|};i)\Big{)}$
where $D(C(bw)\backslash{p_{i}})$ describes population diversity without individual $p_{i}$. For maximising the diversity of blocked edges, we minimise the $SortedD(C(bw)\backslash{p_{i}})$ in lexicographic order, where sorting is performed in descending order.
$\text{{Sorted}}D(C(bw)\backslash{p_{i}})=sort\Big{(}\big{(}c(bw_{1})-(bw_{1};i)\big{)},.......,\big{(}c(bw_{|BW|})-(bw_{|BW|};i)\big{)}\Big{)}$
To maintain diversity, we identify an individual $q$ in the population whose removal maximises overall population diversity, formally, the one that minimises $SortedD(C(bw)\backslash{p_{q}})$. This individual $q$ is removed if it contributes the least to diversity and its fitness score is not near-optimal. However, if the newly created individual has the highest fitness score in the population, we instead remove the individual with the lowest fitness score. Through this selective replacement strategy, the EDO process evolves a diverse and high-quality set of defensive blocking plans, which together define the defender’s policy.
@@ 5.3 Results
Table 2: Attacker success rate on the r1000 AD graph under different approaches (lower values indicate better defensive performance).
[…]
@@ I P N Avg I P N
[…]
Table 3: Attacker success rate on the r2000 AD graph under different approaches (lower values indicate better defensive performance).
[…]
@@ 6 Discussion
Limitations and Future Directions. Our evaluation relies on synthetic AD graphs generated via BloodHound’s DBCreator, which offer structural realism but do not fully capture the operational complexity of production environments, such as evolving user behaviours, dynamic access policies, and diverse edge types. Additionally, structural patterns and attack paths in real-world AD deployments may differ from those in synthetic graphs, which could impact the generalisability of learned defence policies. Future work will aim to validate the framework on production AD environments and incorporate richer edge semantics and behavioural dynamics.

### v2

@@ 1 Introduction
Enterprise networks are increasingly targeted by sophisticated adversaries who exploit configuration flaws and access dependencies to escalate privileges and compromise high-value assets [3, 16]. Defenders, meanwhile, must respond under uncertainty and with limited resources, making proactive, strategic countermeasures essential. These challenges are amplified in complex enterprise infrastructures, where hierarchical access and interdependent components expand the attack surface [13, 14, 7, 4].
A particularly critical target in this landscape is Microsoft Active Directory (AD), the central identity and access management system for Windows-based networks [9, 8, 6]. Over 90% of Fortune 1000 companies rely on AD to govern authentication, authorisation, and access control [15], making it a high-value target. According to Enterprise Management Associates (EMA 2021), more than half of surveyed organisations experienced AD-related breaches. These breaches often follow an “identity snowball” pattern, where attackers pivot from low-privilege accounts toward the highly privileged Domain Admin (DA) node via lateral movement [5]. AD environments can be represented as attack graphs, with nodes denoting entities (e.g., users, machines, groups) and edges representing access relationships. Tools such as BloodHound [2] are used to analyse AD graphs by identifying shortest paths to high-value targets, but they assume static attacker strategies (these strategies assume fixed, precomputed paths that remain unchanged despite defensive actions) and fail to model adaptive behaviour. In practice, attackers frequently adapt to blocked or failed paths by exploiting alternate paths, credentials, underused privileges, and overlooked misconfigurations. Prior work for defending AD graphs has explored edge-blocking techniques that selectively block high-impact edges. However, these methods often assume fixed attacker strategies and lack mechanisms to model dynamic adaptation, limiting their robustness and generalisability to real-world threats.
@@ 4.1 Graph Reduction via Fixed-Parameter Tractable Procedure
Condensing the AD Graph. We propose a Fixed-Parameter Tractable (FPT) procedure to preprocess the original AD graph into a smaller yet strategically equivalent representation. Intuitively, an AD graph can be viewed as a spanning tree augmented by $h$ feedback edges, where h=m(n1)h=m-(n-1) and $m$ and $n$ denote the number of edges and nodes, respectively. Within this graph, Splitting nodes ($t$) are nodes with multiple outgoing edges, representing critical decision points for the attacker, and Entry nodes ($s$) are nodes from which the attacker can initiate an attack. Let SplitSet represent the set of all splitting nodes and EntrySet represent the set of all entry nodes. Our FPT procedure leverages these sets to systematically condense the graph while preserving the structural properties critical to attacker–defender interactions. To facilitate this, we use the concept of Non-Splitting Paths [10]:
@@
By condensing each NSP into a single super-edge, our FPT procedure reduces the attacker’s overall state space from $3^{|Edges|}$ to $3^{|\text{NSP}|}$. The resulting condensed graph contains ($|\textsc{SplitSet}|+|\textsc{EntrySet}|+1$ DA) nodes and $|\text{NSP}|$ edges. This reduction preserves the critical connectivity and decision points of the original graph while significantly improving tractability for algorithmic design.
Modelling the Attacker’s Problem as a Markov Decision Process. We formulate the attacker’s decision-making process as a Markov Decision Process (MDP), wherein each state $s$ is represented as a vector of length $|\text{NSP}|$. Each element of the state vector encodes the status of a corresponding NSP and takes one of three values: ‘F’, ‘S’, ‘?’. The state vector can be represented by:
$\text{State vector s:}\quad\underbrace{<F,?,?,?,F,?,?,S,?,?,S,F>}_{\text{Length = $|\text{NSP}|$}}$ (1)
@@ 4.3 Defender Strategy Optimisation via Evolutionary Diversity Optimisation (EDO)
$\text{Defensive vector:}\quad\langle 0,1,0,\ldots,1,0,1\rangle$ (6)
@@
The newly generated individual is retained in the population if its fitness, defined as the attacker’s success rate under the corresponding blocking plan, as predicted by the GNN, falls within a tolerance window around the current best score. Specifically, it must lie within the range $f^{*}-0.1,\ f^{*}+0.1]$, where $f^{*}$ is the best fitness value observed so far. This 0.1 margin ensures that near-optimal plans are preserved while maintaining population diversity, unless the new individual significantly improves overall diversity.
@@
[…]
@@ 5.3 Results
Table 2: Comparison of attacker success rates on the r1000 Active Directory graph across different defence strategies. Lower values indicate stronger defensive effectiveness.
  Approach Success Rate (%) Time (hours)
@@ I P N Avg I P N
Table 3: Comparison of attacker success rates on the r2000 Active Directory graph across different defence strategies. Lower values indicate stronger defensive effectiveness.
  Approach Success Rate (%) Time (hours)
@@ 6 Discussion
Limitations and Future Directions. Our evaluation relies on synthetic AD graphs generated via BloodHound’s DBCreator, which offer structural realism but do not fully capture the operational complexity of production environments, such as evolving user behaviours, dynamic access policies, and diverse edge types. Additionally, structural patterns and attack paths in real-world AD deployments may differ from those in synthetic graphs, which could impact the generalisability of learned defence policies. Future work will aim to validate the framework on production AD environments, incorporating richer edge semantics and behavioural dynamics, and explore the integration of Large Concept Models (LCMs) [1] as high-level reasoning layers to augment graph-based learning with contextual understanding of vulnerabilities, adaptive defence narratives, and automated policy explanations.